What are the EXACT values of the roots of the quadratic equation 1/2(x^2-6x)=2 ?
Answer by stanbon(75887)

What are the EXACT values of the roots of the quadratic equation
1/2(x^2-6x)=2 ?
-------------
x^2-6x = 4
x^2-6x-4 = 0
x = [6 +- sqrt(36-4*-4)]/2
x = [6 +- sqrt(52)]/2
x = [6 +- 2sqrt(13)]/2
x = 3+sqrt(13) or x = 3-sqrt(13)
=====================
Cheers,
Stan H.